Calculate the height of a satellite from the surface of the earth if its period is 12 h. Compare this height with the radius of the earth.
1
Expert's answer
2020-03-06T10:26:48-0500

"m\frac{v^2}{r}=m\frac{GM}{r^2}"

"v=\frac{2\pi r}{T}"

So,


"r^3=(R+h)^3=\frac{GM}{4\pi^2}T^2"

"(6.38\cdot 10^{6}+h)^3=\frac{(6.67\cdot 10^{-11})(5.97\cdot 10^{24})}{4\pi^2}(12\cdot 3600)^2"

"h=20.22\cdot 10^{6}\ m"

"\frac{h}{R}=\frac{20.22\cdot 10^{6}}{6.38\cdot 10^{6}}=3.17"